With a mass of 25 g hung from its end, the total length of a spring is 15 cm. When the mass is increased to 40 g, the total length of the spring is 18 cm. The unstretched length of the spring is
1
Expert's answer
2020-06-30T18:24:22-0400
"mg=k(l-x)\\Mg=k(L-x)\\\frac{m}{M}=\frac{l-x}{L-x}\\\frac{25}{40}=\frac{15-x}{18-x}\\x=10\ cm"